OpenStudy (anonymous):

A factor of 6x^2 + 8x^2 -30y - 40 would be ... 3y + 4?

OpenStudy (anonymous):

Do you mean this? \[6x^2 + 8x^2 -30x - 40\] If 3x+4 is a factor, then x=-4/3 would be a root. \[6x^2 + 8x^2 -30x - 40=0\] Plug x=-4/3 and tell me what do you get.